OpenStudy (anonymous):

The numerator of a fraction is 3 times the denominator. If 5 is subtracted from the numerator, and added to the denominator, the resulting fraction is equal to 16/12

OpenStudy (anonymous):

x = 7

OpenStudy (anonymous):

Solve the following for x:\[\frac{3x-5}{x+5}=\frac{16}{12} \]